OpenStudy (anonymous):

If you are finding the line parallel to the graph of x + 3y = 5, what is the slope of the line?

OpenStudy (anonymous):

when y= mx+ c is a line, m is the slope. here m is -1/3. which is the slope. for starters, reorder terms so that it matches the order y= mx+ c

OpenStudy (anonymous):

-3! is the asnwer!!

OpenStudy (anonymous):

oh sorry..-1/3!!

OpenStudy (anonymous):

parallel lines have the same slope: thus it will have m=-1/3

OpenStudy (anonymous):

x + 3y = 5 x + 3y -x = 5 - x 3y = -x + 5 3y/3 = (-x+5)/3 y = -1/3x + 5/3 so m = -1/3 here's the explanation in case u need it!!